OpenStudy (anonymous):

Complete the equations of the system in slope-intercept form. Use a decimal for the slope if necessary. Line 1 x y 0 3 1 5 Line 2 x y –1 1 –2 –1 Line 1: y = x + Line 2: y = x +

OpenStudy (anonymous):

@DannyO19

OpenStudy (anonymous):

@BlueMiku

OpenStudy (bluemiku):

lol how did you find me

OpenStudy (anonymous):

me friend said to tag you

OpenStudy (bluemiku):

what really? who?

OpenStudy (anonymous):

idk what their name is on here

RhondaSommer (rhondasommer):

slope intercept form is mx+b

RhondaSommer (rhondasommer):

y=mx+b

RhondaSommer (rhondasommer):

second y coodinate - first y coordinate m=--------------------------------- second x coodinate - first x coordinate

OpenStudy (anonymous):

could you explain it a little better

pooja195 (pooja195):

What she means is the slope formula \[\huge~\rm~\bf~ \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (anonymous):

so you would take for line 1 5-3 over 1-0

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

Isn't the coordinates (0,3) (1,5) ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

y = o.5x+1

OpenStudy (anonymous):

I think that's line 1

OpenStudy (anonymous):

I haven't done this in 3 years...

OpenStudy (anonymous):

4*

OpenStudy (tootzrll):

okay, so line 1 has these x and y coordinates: (0,3) and (1,5) line 2 has these x and y coordinates: (-1,1) and (-2,-1) (0,3) and (1,5) are most likely to be on the same line, hence "line 1". same with the coordinates for line 2.

OpenStudy (tootzrll):

what u need to do is plot line 1 first, then draw the lines and connect the two point to each line. then do the same with line 2

OpenStudy (anonymous):

toot is a brainiac

OpenStudy (tootzrll):

xD no just good with graphs

OpenStudy (mathmale):

@michelleball477: 1. What is the slope of the line in question? 2. What is the general point-slope equation for a straight line? 3. What is the point-slope equation of this particular straight line?